Use this information about Department & to answer the question that folow Department G had 1.920 units 2ร% completed at t g ofn-pernd, 12,200 unts were completed drng the perod,L600units were 20% completed at the erdofdw perod, and th·11н﹃ muda turge es- debited to the departmental work m pracess account during the penod Work in process, beginning of period Couts added during penod Direct materials (31,880 unts at $9 Direct labor Factory everhead 26,000 06,920 %,120 6,120 Oa $26,000 b.$38,672 c. $41,840 Od. $43,440
Answer #1

Under First in first method ,units started first are transfered out first .

out of 12200 units completed during the period ,1920 units are from beginning WIP and remaining (12200-1920) =10280 units are from units started and complete during the period .

Units Direct Material Direct labor Factory overhead
Beginning WIP 0 1920*75% = 1440    [25%already completed] 1440
units started and completed 10280 10280 10280
Ending WIP 1600 1600*20%=320 320
Equivalent unit 11880 12040 12040
cost incurred during the period 106920 96320 36120
cost per equivalent unit

106920/11880

=$ 9

96320/12040= $ 8

36120/12040

=$ 3

Total cost of Beginning WIP completed during the period =beginning cost + current cost [0*9]+[1440*8]+[1440*3]

               =26000+[ 0+11520+4320]

                = 26000+15840

              = 41840

correct option is "C"